The random variable J has a geometric distribution and it is given that, ๐‘ƒ(๐ฝ=5)/๐‘ƒ(๐ฝ=7) = 25/16 .


Find ๐‘ƒ(๐ฝ = 4)


1
Expert's answer
2021-01-12T01:05:42-0500

For the geometric distribution



"P(J=j) = p(1-p)^{j-1}"

Therefore



"\frac{P(J=5)}{P(J=7)} = \frac{p(1-p)^{5-1}}{p(1-p)^{7-1}}=\frac{25}{16}"


"\frac{(1-p)^{4}}{(1-p)^{6}}=\frac{25}{16}""(1-p)^{-2}=\frac{25}{16}""-2 * log (1-p)=log (\frac{25}{16})""-2 * log (1-p)=0.1938""log (1-p)=-0.0969""1-p = e^{-0.0969}""p=1-0.90763767804293""p=0.0924"

The PDF is therefore


"P(J=j) = 0.0924(0.9076)^{j-1}""P(J=4) = 0.0924(0.9076)^{4-1}""P(J=4) =0.0691"